Assume today’s settlement price on a CME EUR futures contract is $1.3240/EUR. You have a long position in one contract.

1.)How do you Calculate the initial performance bond account balance?

2.)Assume that the next three days’ settlement prices are $1.3226, 1.3233, and 1.3249.

Calculate the changes in the performance bond account from daily marking-to-market and

the balance of the performance bond account after the third day.
Answer #1

Ans

Note:

The data in the question in insufficient to solve part 1 because performance account current balance is not mentioned. However i will solve for the other part. In case some data is required i will specify variable just place the value of that variable to get the final answer.

Solution

Assume that today’s settlement price on a CME EUR futures contract is $1.3240/EUR.

Here we have long position in one contract. This means that whenever the price will rise our mark to market will be positive , that means we will gain for that day. And when the price will decrease as compared to previous day we will post a loss, as we are in long position.

Your performance bond account currently has a balance of $X.

The next three days settlement prices are as follows: $1.3226, $1.3233, and $1.3249. Calculate the changes in the performance account from daily marking to market.

Your initial account balance is $X and you have taken a long position on Y units of Euro at a price of $1.3240.

Here the variable $X initial performance and units of Euro should be given, in the question posted its not mentioned. Kindly check and replace the value with the variable to get appropriate answer.

Performance bond account from Daily Mark to Market


Day 1:

Closing price = $1.3226.

You post a LOSS of ($1.3240 – $1.3226) = $0.0014. per unit of EURO

Therefore total loss will be A = 0.0014 * (Y units of Euro )

This is subtracted from your account. Your current balance is $X - $A = $X1.----------------------------(1)


Day 2:

Closing price = $1.3233.

You post a PROFIT of ($1.3226 – $1.3233) = $0.0007.

Therefore total PROFIT will be $B = 0.0007 * (Y units of Euro )

This is added to your account. Your current balance is $X1 + $B = $X2 -----------------------------(2)


Day 3:

Closing price = $1.3249.

You post a PROFIT of ($1.3233 – $1.3249) = $0.0016.

Therefore total PROFIT will be $C = 0.0016 * (Y units of Euro )

This is added to your account. Your current balance is $X2 + $C = $X3 --------------------------------(3)

The closing balance of the performance bond account after the third day will be $X3

Note: cross-check the question for value of initial bond performance ($X) and units of Euro for long position (Y) . Replace the value for variables and you will get $X3 .